### v2

@@ 1 Introduction
Neural synchronization between distinct brain regions underlies critical cognitive processes including perception, attention, and memory formation [4],[10]. Disruptions in inter-regional synchronization are implicated in neurological disorders such as epilepsy, Alzheimer’s disease, and schizophrenia [9]. Despite its significance, the structural determinants of synchronization propagation, particularly the role of nodal centrality and inter-areal connectivity, remain incompletely characterized.
@@ 2.1 Network Architecture and Topology Modification
Neural populations were modeled as undirected stochastic block model (SBM) networks [7] with two equally sized clusters ($N=500$ neurons, $|V_{t}|=|V_{o}|=250$, where $|V_{t}|$ and $|V_{o}|$ are the sizes of the stimulated and measured populations of neurons). The adjacency matrix $A$ was constructed with intra-cluster connection probability $p_{\mathrm{intra}}=0.15$ and inter-cluster probability $p_{\mathrm{inter}}$ systematically varied from 0.01 to 0.1. The signal was applied to neurons within the top 10%, 15%, or 20% centrality percentiles ($p_{\mathrm{input}}$) or to 10–20% of randomly selected neurons within the cluster $V_{t}$ (Proxy stimulation strategy and Random baseline). To overcome SBM’s inherent homogeneity and in order to introduce topological diversity, we implemented a centrality boosting procedure defined by:
@@
  • $\beta=1.5$ (50% centrality boost)
@@ 2.2 Centrality Measure Computation
We evaluated six topological centrality measures that quantify distinct aspects of node influence in networks. For graph $G=(V,E)$ with adjacency matrix $A$ ($A_{uv}=1$ if edge exists, 0 otherwise), $n=|V|$ nodes, and shortest path distance $d(u,v)$:
@@ 2.3 Neural Dynamics Model
We consider spiking neural network consisting of leaky integrate-and-fire (LIF) neurons [1]. As was mentioned in the previous section, the whole network was divided into two equal populations. The connections between individual neurons within and between populations were established according to the desired adjacency matrix.
Each of the two populations contained $80\%$ excitatory and $20\%$ inhibitory neurons. The equations for the membrane potential of a single neuron were written as follows:
@@
To simulate background neural activity, all neurons received independent stimuli $v_{n}$ with a Poisson distribution. Each spike from the Poisson generator causes an instantaneous increase in membrane potential, which can be represented as the sum of delta functions:
@@
where $t_{n}$ is the time of arrival of the spikes. We assumed that
@@
[…]
2.4 Numerical Experiments and Analysis
Simulation Protocol
  1. 1.
    Network generation (SBM with specified $p_{\mathrm{intra}}$, $p_{\mathrm{inter}}$, 20 times for each parameters pair)
  2. 2.
    Centrality computation and node selection
  3. 3.
    Topology modification ($\Delta k$ edges added)
  4. 4.
    Density compensation (random edge removal)
  5. 5.
[…]
  6. 6.
    100 ms warm-up period excluded from analysis
@@ 3 Results
We ran several experiments, in which we choose specific driver nodes in population $V_{t}$ and measured neural activity in population $V_{o}$. We tested several experimental set-ups:
  • network with and without boosting;
  • selecting driver neurons within the population $V_{t}$ among the nodes with maximum centrality measures;
  • random driver neurons of cluster $V_{t}$ that are also the neighbors of the most central neurons of the cluster $V_{o}$;
  • random driver neurons of cluster $V_{t}$ or random baseline.
These topological control strategies demonstrated different efficiencies. In the absence of centrality boosting, driver node selection strategy significantly impacted the activity propagation efficiency. Crucially, stimulation of top-centrality nodes substantially outperformed proxy stimulation across all parameter set elements (see Fig. 2).
Figure 2: Propagation efficiency in non-boosted networks. This figure depicts the distribution of average activity frequencies in the second network cluster following targeted stimulation of high-centrality nodes within the first cluster, specifically focusing on the top $20\%$ percentile for each centrality metric. Six subgraphs are presented (left to right, top to bottom), corresponding to the following centrality measures: betweenness, closeness, degree, eigenvector, harmonic, and percolation. Each subgraph displays three data planes representing distinct stimulation conditions: the average activity response in the second cluster when stimulating the top-centrality nodes ($V_{t}$) within the first cluster; the response when stimulating proxy nodes ($V_{t}$) in the first cluster; and a random baseline response from stimulating randomly selected node subsets ($10\%,15\%,20\%$) in the first cluster. The top-centrality and proxy stimulation results share a single colormap scale but are visually distinguished by differential transparency, while the random baseline employs a separate colormap scale as visually indicated.
Network Control Dynamics in Structural Connectomes: Non-boosted Network Performance
Our analysis reveals fundamental constraints in stimulus propagation through structural connectomes. Below the critical percolation threshold ($p_{\mathrm{inter}}<0.04$), no strategy elicited significant propagation – indicating a fundamental physical constraint on information transmission in sparse connectomes. Beyond this threshold, strategy-dependent effects emerged:
@@
    Threshold modulation: Boosted networks exhibited significant propagation amplification, compared to non-boosted case 2.3$\times$ increase in propagation efficiency compared to non-boosted networks at $p_{\mathrm{inter}}=0.10$, with the mean firing rate in $V_{o}$ rising from $3.33\pm 0.98$ Hz to $7.69\pm 0.62$ Hz under optimal driver selection. This demonstrates that targeted neuroplasticity can reconfigure information routing pathways to overcome physical connectivity constraints.
@@
    • At $p_{\mathrm{inter}}=0.07$, closeness-centrality targeting dominated ($2.06\pm 0.44\leavevmode\nobreak\ \text{Hz}$), while proxy stimulation remained ineffective ($0.41\pm 0.38\leavevmode\nobreak\ \text{Hz}$, 5.02-fold reduction)
    • At $p_{\mathrm{inter}}=0.10$, closeness-centrality again prevailed ($7.69\pm 0.62\leavevmode\nobreak\ \text{Hz}$), but proxy stimulation showed relative improvement ($2.59\pm 1.48\leavevmode\nobreak\ \text{Hz}$, 2.96-fold reduction)
Figure 3: Propagation efficiency in boosted networks. This figure depicts the distribution of average activity frequencies in the second network cluster following targeted stimulation of high-centrality nodes within the first cluster, specifically focusing on the top $20\%$ percentile for each centrality metric. Six subgraphs are presented (left to right, top to bottom), corresponding to the following centrality measures: betweenness, closeness, degree, eigenvector, harmonic, and percolation. Each subgraph displays three data planes representing distinct stimulation conditions: the average activity response in the second cluster when stimulating the top-centrality nodes ($V_{t}$) within the first cluster; the response when stimulating proxy nodes ($V_{t}$) in the first cluster; and a random baseline response from stimulating randomly selected node subsets ($10\%,15\%,20\%$) in the first cluster. The top-centrality and proxy stimulation results share a single colormap scale but are visually distinguished by differential transparency, while the random baseline employs a separate colormap scale as visually indicated.
Interpretive Synthesis The persistent >65% efficacy gap in proxy stimulation – even in boosted networks – underscores that spatial proximity to hubs cannot compensate for the absence of topological position. Optimal targets consistently exhibited high closeness centrality, reflecting their capacity to integrate information through short functional paths. By contrast, the modest improvement in proxy efficacy under high connectivity with boosting suggests that enhanced network integration partially compensates for suboptimal target selection.
Figure 4: Dynamic propagation characteristics. (A) Non-boosted networks: Top-centrality shows narrowband enhancement at driving frequency (10 Hz). (B) Boosted networks: All strategies exhibit spectral peaks but with varying signal-to-noise ratios (SNR: top=7.2 dB, hybrid=5.1 dB, proxy=2.8 dB).
@@ 4 Discussion
Our results demonstrate that topological centrality provides an effective criterion for driver node selection in modular neural networks. The superior performance of top-centrality nodes stems from their dual advantage: (1) high integration within source populations enables efficient input amplification, and (2) preferential connection to inter-module bridges facilitates propagation.
The limited efficacy of proxy stimulation reveals that spatial proximity to target hubs is insufficient for effective control, aligning with recent findings that optimal control requires both local integration and global bridging capabilities [5]. Spectral analysis confirms that centrality-based selection preserves temporal fidelity of transmitted signals.
Notably, the performance gap between selection strategies was most pronounced at critical $p_{\mathrm{inter}}$ values (0.07), corresponding to theoretical percolation thresholds for modular networks. This indicates topological control strategies are particularly valuable near network phase transitions—precisely where biological neural systems typically operate.
@@ 5 Conclusion
  • Metric-specific performance: Degree and closeness centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement (>5×>5$\times$> 5 × advantage over proxy strategies)

### v3

@@ 1 Introduction
Neural synchronization between distinct brain regions underlies critical cognitive processes, including perception, attention, and memory formation [1],[2]. Disruptions in inter-regional synchronization are implicated in neurological disorders such as epilepsy, Alzheimer’s disease, and schizophrenia [3]. Despite its significance, the structural determinants of synchronization propagation, particularly the role of nodal centrality and inter-areal connectivity, remain incompletely characterized.
@@ 2.1 Network Architecture and Topology Modification
Neural populations were modeled as undirected stochastic block model (SBM) networks [5] with two equally sized clusters ($N=500$ is the total number of neurons, $|V_{t}|=|V_{o}|=250$, where $|V_{t}|$ and $|V_{o}|$ are the sizes of the stimulated $V_{t}$ and measured $V_{o}$ populations of neurons). The adjacency matrix $A$ was constructed with intra-cluster connection probability $p_{\mathrm{intra}}=0.15$ and inter-cluster probability $p_{\mathrm{inter}}$ systematically varied from 0.01 to 0.1. The signal was applied to neurons within the top 10%, 15%, or 20% centrality percentiles ($p_{\mathrm{input}}$) or to 10–20% of randomly selected neurons within the cluster $V_{t}$ (Proxy stimulation strategy and Random baseline). To overcome SBM’s inherent homogeneity and in order to introduce topological diversity, we implemented a centrality boosting procedure defined by:
@@
  • $S$ is the set of selected driver neurons
@@ 2.2 Centrality Measure Computation
We evaluated six topological centrality measures that quantify distinct aspects of node influence in networks. For graph $G=(V,E)$ with adjacency matrix $A$ ($A_{uv}=1$ if edge exists, 0 otherwise), $n=|V|$ nodes, and shortest path distance $d(u,v)$ we define the following centrality measures:
@@ 2.3 Neural Dynamics Model
We consider spiking neural network consisting of leaky integrate-and-fire (LIF) neurons [8]. As was mentioned in the previous section, the whole network was divided into two equal populations. The connections between individual neurons within and between populations were established according to the desired adjacency matrix. Each of the two populations contained $80\%$ excitatory and $20\%$ inhibitory neurons.
The equations for the membrane potential of a single neuron were written as follows:
@@
[…]
@@
where $t_{n}$ is the time of arrival of the spikes. We set
@@
2.4 Numerical Experiments
We conducted several experiments in which we chose specific driver nodes in the population $V_{t}$ (with a fraction $p_{input}$ of the total amount) and measured neural activity in the population $V_{o}$. Several experimental protocols were implemented. For the first set of experiments, the SBM network was generated with the specified $p_{\mathrm{intra}}$, $p_{\mathrm{inter}}$. Next, the centrality measures were computed for the nodes of the $V_{t}$ population. 10-to-20% of the nodes with the highest centralities in the population $V_{t}$ were chosen as driver nodes. The dynamics of the resulting network was simulated for 5 seconds of biological time in Brian 2 neural simulator [9]. The initial 100 ms warm-up period was excluded from the analysis.
In an alternative setting, the centrality measures were computed for the nodes of the $V_{o}$ population. The driver nodes were randomly chosen among the nodes of $V_{t}$ that are neighbors of the top central nodes in $V_{o}$.
The above experimental settings were compared to the baseline experiment in which the driver nodes were chosen in the population $V_{t}$ randomly. The simulations were repeated 20 times for each set of parameters. The results of simulations are shown in Fig. 1.
Figure 1: Propagation efficiency in non-boosted networks. This figure depicts the distribution of average activity in the $V_{o}$ population following targeted stimulation of high-centrality nodes within the $V_{t}$ population. Six subgraphs are presented (left to right, top to bottom), corresponding to the following centrality measures: betweenness, closeness, degree, eigenvector, harmonic, and percolation. Each subgraph displays three surfaces representing distinct stimulation conditions: the average activity response in the $V_{o}$ cluster when stimulating the top-centrality nodes within the $V_{t}$ cluster; the response when stimulating proxy nodes in the $V_{t}$ cluster; and a random baseline response from stimulating randomly selected driver node subsets ($p_{input}=10\%,15\%,20\%$) in the $V_{t}$ cluster. The top-centrality and proxy stimulation results share a single colormap scale but are visually distinguished by differential transparency, while the random baseline employs a separate colormap scale as visually indicated
In the second set of experiments, the experimental procedure was similar to the previous one except the additional boosting applied to the $n_{boost}$ nodes with the highest centrality (Fig. 2):
  • Topology of the network was modified by adding $\Delta k$ edges.
  • The increased density of edges was compensated by random edge removal.
The results of these simulations are shown in Fig. 3.
Figure 3: Propagation efficiency in boosted networks. This figure depicts the distribution of average activity in the $V_{o}$ population following targeted stimulation of high-centrality nodes within the $V_{t}$ population. Six subgraphs are presented (left to right, top to bottom), corresponding to the following centrality measures: betweenness, closeness, degree, eigenvector, harmonic, and percolation. Each subgraph displays three surfaces representing distinct stimulation conditions: the average activity response in the $V_{o}$ cluster when stimulating the top-centrality nodes within the $V_{t}$ cluster; the response when stimulating proxy nodes in the $V_{t}$ cluster; and a random baseline response from stimulating randomly selected node subsets ($p_{input}=10\%,15\%,20\%$) in the $V_{t}$ cluster. The top-centrality and proxy stimulation results share a single colormap scale but are visually distinguished by differential transparency, while the random baseline employs a separate colormap scale as visually indicated
@@ 3 Results
The control strategies described in the previous section demonstrated different efficiencies in activity propagation between populations. In the absence of centrality boosting, the driver nodes selection strategy significantly impacted the activity propagation efficiency. Crucially, stimulation of the top centrality nodes substantially outperformed proxy stimulation across all parameter set elements (see Fig. 1).
In non-boosted networks, our analysis reveals fundamental constraints in stimulus propagation through structural connectomes. Below the critical percolation threshold ($p_{\mathrm{inter}}<0.04$), no strategy elicited significant propagation – indicating a fundamental physical constraint on information transmission in sparse connectomes. Beyond this threshold, strategy-dependent effects emerged:
@@
    Threshold modulation: Boosted networks exhibited significant propagation amplification, compared to non-boosted case 3.15$\times$ increase in propagation efficiency compared to non-boosted networks at $p_{\mathrm{inter}}=0.10$, with the mean firing rate in $V_{o}$ rising from $3.33\pm 0.98$ Hz to $10.5\pm 0.13$ Hz under optimal driver selection. This demonstrates that targeted neuroplasticity can reconfigure information routing pathways to overcome physical connectivity constraints.
@@
    • At $p_{\mathrm{inter}}=0.07$, closeness-centrality targeting dominated ($7\pm 0.69\leavevmode\nobreak\ \text{Hz}$), while proxy stimulation remained ineffective ($1.8\pm 1.23\leavevmode\nobreak\ \text{Hz}$, 3.89-fold reduction)
    • At $p_{\mathrm{inter}}=0.10$, closeness-centrality again prevailed ($10.5\pm 0.13\leavevmode\nobreak\ \text{Hz}$), but proxy stimulation showed relative improvement ($4.7\pm 2.07\leavevmode\nobreak\ \text{Hz}$, 2.23-fold reduction)
Interpretive Synthesis The persistent >55% efficacy gap in proxy stimulation – even in boosted networks – underscores that spatial proximity to hubs cannot compensate for the absence of topological position. Optimal targets consistently exhibited high closeness centrality, reflecting their capacity to integrate information through short functional paths. By contrast, the modest improvement in proxy efficacy under high connectivity with boosting suggests that enhanced network integration partially compensates for suboptimal target selection.
Figure 4: Dynamic propagation characteristics. (A) Non-boosted networks: Top-centrality shows narrowband enhancement at driving frequency (10 Hz). (B) Boosted networks: All strategies exhibit spectral peaks but with varying signal-to-noise ratios (SNR: top=8.27 dB, proxy=7.49 dB)
@@ 4 Discussion
Our results demonstrate that topological centrality provides an effective criterion for driver nodes selection in modular neural networks. The superior performance of top-centrality nodes stems from their dual advantage: (1) high integration within source populations enables efficient input amplification, and (2) preferential connection to inter-module bridges facilitates propagation.
The limited efficacy of proxy stimulation reveals that spatial proximity to target hubs is insufficient for effective control, aligning with recent findings that optimal control requires both local integration and global bridging capabilities [10]. Spectral analysis confirms that centrality-based selection preserves the temporal fidelity of transmitted signals (Fig. 4).
Notably, the performance gap between selection strategies was most pronounced at critical $p_{\mathrm{inter}}$ values (0.07), corresponding to theoretical percolation thresholds for modular networks. This indicates topological control strategies are particularly valuable near network phase transitions, precisely where biological neural systems typically operate.
@@ 5 Conclusion
  • Metric-specific performance: Closeness and harmonic centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement (>5×>5$\times$> 5 × advantage over proxy strategies)